# Visual Scratchpads: Enabling Global Reasoning in Vision

### v1

@@ D.1 Additional model scaling experiments on maze (circular)
In the model scaling experiments conducted on the maze circular dataset in Figure 9, we observe a similar behavior to that seen on maze rectangular. For larger model sizes (Base, Large, and Huge), both the inductive and single-frame scratchpads achieve near-perfect accuracy. However, the inductive scratchpad particularly shines when it comes to smaller models. With the ViT Small model, the inductive approach significantly outperforms the single-frame scratchpad, yielding a performance improvement of more than 30 percentage points. This indicates the effectiveness of the inductive method in handling resource-constrained settings, maintaining superior OOD generalization even when model capacity is limited.
@@
Figure 9: Maze (circular) 16 model size experiments. The model behavior is similar to the maze (rectangular) dataset. Inductive scratchpad is on par with Single-frame for B, L and H but has a significant advantage on S.
D.2 Relationship between model size and OOD generalization
@@
The plot in Figure 10 presents the OOD generalization performance for models of different sizes (B and H) trained on task complexity 12 and tested on more complex tasks ranging from 14 to 24. Notably, the inductive scratchpad consistently outperforms the single-frame scratchpad across the entire range of task complexities, irrespective of model size. This trend holds true for both the base and huge models, although the performance gap between the two approaches seems to decrease as model size increases. This suggests that the single-frame scratchpad can somewhat benefit from larger models. However, as shown in the main paper, a key advantage of the inductive scratchpad lies in its ability to improve performance by expending more compute at inference time, enabling smaller models to perform well. We hypothesize that the diminishing gap in performance with the huge model might be attributed to it being more data-hungry. Since the inductive scratchpad sees two steps per iteration for each sample, it may be more prone to memorization, suffering from the additional exposure to images during training.
Figure 10: OOD experiment with model size scaling
D.3 Additional OOD experiments on maze (circular) and strings datasets
@@
Similar to the experiments presented in the main paper, on the maze circular dataset (see Table 4), both the inductive and single-frame scratchpads achieve near-perfect performance on in-distribution (ID) tasks. However, for out-of-distribution (OOD) tasks, the inductive scratchpad significantly outperforms the single-frame scratchpad, achieving 96.88% accuracy compared to 62.99%. This trend mirrors the results observed on the maze rectangular dataset, where OOD generalization is again much better for the inductive method. For the strings dataset (see Figure 11), the pattern slightly differs. Strings is a more challenging dataset overall, as established in the main paper, which makes OOD generalization particularly difficult. Nonetheless, the inductive scratchpad consistently performs better than the single-frame scratchpad, especially on more complex OOD tasks, with the exception of size 14, which is the simplest OOD task in this setting.
@@
Single-frame scratchpad 100.00 62.99
Inductive scratchpad 99.98 96.88
@@
Table 4: In-distribution (ID) and out-of-distribution (OOD) performance on the maze circular dataset for different methods. The inductive scratchpad outperforms the single-frame scratchpad in the OOD setting.
D.4 Additional ablations for the multi-frame baseline
@@
In the main paper, we discussed the factors contributing to the success of the inductive scratchpad model, including increased supervision during training, the halting mechanism, and the integration of teacher forcing with training on the output distribution. In this section, we present additional experiments to evaluate the impact of the multi-frame supervision. We introduced a multi-frame baseline model, which, while similar to the single-frame model, features multiple heads for predicting several scratchpad frames.
@@
Our previous findings indicated that the multi-frame baseline did not yield any performance gains on OOD samples compared to the single-frame model. However, there is a scenario where the multi-frame approach proves beneficial: it aids convergence for smaller models (Base and Large). As illustrated in Figure 12, the inductive scratchpad converges across all model sizes (Small, Base, Large, and Huge), while the single-frame scratchpad only converges for the Huge model, indicating a greater computational demand to find the solution. The multi-frame model mitigates this issue by facilitating convergence in Base and Large models, suggesting that while it may still struggle with OOD, as noted in the main paper, it may help model in discovering the solution. This improvement can be attributed to the presence of additional frames, which provide better guidance on the path to reaching the solution.
Figure 12: Scaling parameters, Single-frame vs. Multi-frame vs Inductive scratchpad.
D.5 Additional staircase examples
While the main paper presents enhanced visualizations of the staircase phenomenon for clarity, it’s important to note that this behavior is also evident in the non-enhanced outputs. As shown in Figure 13, the third row displays the raw model outputs for the cycles 16 task, which exhibit the same progressive learning pattern described earlier. This confirms that the staircase effect is not an artifact of post-processing but a genuine characteristic of the model’s learning process.
Moreover, this hierarchical learning phenomenon is not limited to the cycles task. Figure 14 demonstrates that a similar staircase behavior emerges in the more complex maze (rectangular) task. In this case, the model’s behavior resembles a spreading “cloud” that progressively discovers contiguous areas of the maze. This is particularly noteworthy because the model is trained only on the final, fully solved maze configuration (shown in the second row of each column), with the first row representing the input maze.
In both the cycles and maze tasks, we observe a consistent pattern of the model first solving easier, more local aspects of the problem before progressively tackling more global structures. This aligns with the theoretical understanding of the staircase effect in learning sparse Boolean functions, but now demonstrated in more diverse and visually complex domains.
Figure 13: Expanded staircase examples for the cycles 16 task.
D.6 Additional scratchpad examples for other tasks
@@
This subsection provides example scratchpads for several tasks, demonstrating target frames for the model. The following figures illustrate scratchpads for tasks like connected and disconnected cycles, strings, and solvable and non-solvable mazes. In Figure 15, we show examples of the cycles 20 dataset with connected cycles. In Figures 16 and 17, scratchpads for the strings 12 dataset with disconnected strings are shown. For maze tasks, Figures 18 and 19 display scratchpads for solvable and non-solvable rectangular mazes. Finally, Figures 20 and 21, do the same for maze circular 16.
Figure 14: Additional staircase examples for the maze (rectangular) task.
Figure 15: Example of scratchpads for the cycles 20 dataset, connected cycles.
Figure 16: Example of scratchpads for the strings 12 dataset, disconnected strings.
Figure 17: Example of scratchpads for the strings 12 dataset, connected strings.
Figure 18: Example of scratchpads for the maze (rectangular) 24 dataset, non-solvable maze.
Figure 19: Example of scratchpads for the maze (rectangular) 24 dataset, solvable maze.
Figure 20: Example of scratchpads for the maze (circular) 16 dataset, non-solvable maze maze.
Figure 21: Example of scratchpads for the maze (circular) 16 dataset, solvable maze.

### v2

@@ D.2 Additional model scaling experiments on maze (circular)
In the model scaling experiments conducted on the maze (circular) dataset in Figure 10, we observe a similar behavior to that seen on maze rectangular. For larger model sizes (Base, Large, and Huge), both the CoS and the single-frame achieve near-perfect accuracy. However, the CoS model particularly shines when it comes to smaller models. With the ViT-Small model, the CoS approach significantly outperforms the single-frame, yielding a performance improvement of more than $30\%$. This indicates the effectiveness of the inductive method in handling resource-constrained settings.
@@
Figure 10: Maze (circular) 16 model size experiments. The model behavior is similar to the maze (rectangular) dataset. CoS is on par with Single-frame for B, L and H but has a significant advantage on S.
D.3 Relationship between model size and OOD generalization
@@
The plot in Figure 11 presents the OOD generalization performance for models of different sizes (B and H) trained on task complexity 12 and tested on more complex tasks ranging from 14 to 24. Notably, the inductive CoS model consistently outperforms the single-frame CoS across the entire range of task complexities, irrespective of model size. This trend holds true for both the base and huge models, although the performance gap between the two approaches seems to decrease as model size increases. This suggests that the single-frame model can somewhat benefit from larger models. However, as shown in the main paper, a key advantage of the CoS lies in its ability to improve performance by using more compute at inference time, enabling smaller models to perform well. We hypothesize that the diminishing gap in performance with the huge model might be attributed to it being more data-hungry. Since the CoS model sees two steps per iteration for each sample, it may be more prone to memorization, suffering from the additional exposure to images during training.
Figure 11: OOD experiments with model size scaling
D.4 Additional OOD experiments on maze (circular) and strings datasets
@@
Similar to the experiments presented in the main paper, on the maze (circular) dataset (see Table 6), both the inductive and single-frame CoS achieve near-perfect performance on in-distribution (ID) tasks. However, for out-of-distribution (OOD) tasks, the inductive CoS model significantly outperforms the single-frame CoS, achieving 96.88% accuracy compared to 62.99%. This trend mirrors the results observed on the maze rectangular dataset, where OOD generalization is again much better for the inductive method. For the strings dataset (see Figure 12), the pattern slightly differs. The strings dataset is a more challenging dataset overall, as established in the main paper, which makes OOD generalization particularly difficult. Nonetheless, the CoS model consistently performs better than the single-frame CoS, especially on more complex OOD tasks, with the exception of size 14, the simplest OOD task in this setting.
@@
single-frame CoS 100.00 62.99
inductive CoS 99.98 96.88
@@
Table 6: In-distribution (ID) and out-of-distribution (OOD) performance on the maze (circular) dataset for different methods. The inductive CoS outperforms the single-frame model in the OOD setting.
D.5 Additional ablations for the multi-frame CoS
@@
In the main paper, we discussed the factors contributing to the success of the inductive CoS model, including increased supervision during training, the halting mechanism, and the integration of teacher forcing with training on the output distribution. In this section, we present additional experiments to evaluate the impact of the multi-frame supervision. We introduced a multi-frame CoS model, which, while similar to the single-frame model, features multiple heads for predicting several CoS frames.
@@
Our previous findings indicated that the multi-frame CoS did not yield any performance gains on OOD samples compared to the single-frame model. However, there is a scenario where the multi-frame approach proves beneficial: it aids convergence for smaller models (Base and Large). As illustrated in Figure 13, the inductive CoS converges across all model sizes (Small, Base, Large, and Huge), while the single-frame CoS only converges for the Huge model, indicating a greater computational demand to find the solution. The multi-frame model mitigates this issue by facilitating convergence in Base and Large models, suggesting that while it may still struggle with OOD, as noted in the main paper, it may help learning in-distribution. This improvement can be attributed to the presence of additional frames, which provide better guidance on the path to reaching the solution.
Figure 13: Scaling parameters, single-frame vs. multi-frame vs. inductive CoS.
Appendix E Globality details
Our main regime of interest, the regime with significant patches (which we studied in the main text), assumes that patches are large enough, e.g., $P\times P$ sized-patch with $P=\sqrt{n}$, such that only a unique ordering of the patches is valid (if one permutes the patches the new sample does not belong to the distribution’s support w.h.p. as in many computer vision tasks of interest). In this case, the histogram part should not be inserted, as is done in our current definition. We used this assumption for defining globality degree in visual domain, which we consider to be the right regime to better understand the targets of interest. In the small-patch regime where the patches have a constant size, i.e., $P\times P$ sized-patch with $P=O_{n}(1)$, one would also need to update slightly the definition of the globality degree in order to capture the fact that targets that are permutation invariant may be more easily learned by the Transformer after dropping the positional embeddings, which results in adding the histogram of the patches to the mutual information as was done in the globality degree for text by Abbe et al. [2024].
Further, note that to define our asymptotic quantities properly, we assume that the number of patches $n$ scales (e.g., as the size of the maze increases, we need a higher resolution image to solve it) [^9] The number of patches in ViT models is usually constant as the images are resized unless the image is so fine-grained that a higher resolution is required, e.g., when the number of graph nodes diverges.. In this case, the requirement would be to have $\alpha=n^{-O_{n}(1)}$ and $k^{*}=O_{n}(1)$ in order for the whole learning complexity to be polynomial in $n$. (Here by learning with polynomial complexity, we mean learning with polynomial model size in polynomial time.)
Appendix F Additional figures
F.1 Chain-of-sketch examples for other tasks
@@
This section provides example chain-of-sketch instances for several tasks, demonstrating target frames for the model. The following figures illustrate chain-of-sketches for tasks like connected and disconnected cycles, strings, and solvable and non-solvable mazes. In Figure 14, we show examples of the cycles 20 dataset with connected cycles. In Figure 15 and 16, sketches for the strings 12 dataset with disconnected and connected strings are shown. For maze tasks, Figures 18 and 17 display sketches for solvable and non-solvable rectangular mazes. Figures 20 and 19, do the same for maze circular 16. Finally, Figure 21 shows the sketches for a PVR task on $7\times 7$ grid and $k=4$ images per row where the parity function is used on the row.
Figure 14: Example of sketches for the cycles 20 dataset, connected cycles.
Figure 15: Example of sketches for the strings 12 dataset, disconnected strings.
Figure 16: Example of sketches for the strings 12 dataset, connected strings.
Figure 17: Example of sketches for the maze (rectangular) 24 dataset, non-solvable maze.
Figure 18: Example of sketches for the maze (rectangular) 24 dataset, solvable maze.
Figure 19: Example of sketches for the maze (circular) 16 dataset, non-solvable maze maze.
Figure 20: Example of sketches for the maze (circular) 16 dataset, solvable maze.
Figure 21: Example of sketches for PVR task on $7\times 7$ grid and $k=4$ images per row where the parity function is used on the row. The final answer which is the parity of the airplane class in the indicated row is zero.
F.2 Additional staircase figures
Figures 22 and 23 present the staircase hierarchical learning phenomenon for cycles and maze datasets. In each figure, the first row presents the input image, the second row presents the ground truth (single-frame) CoS. The third row presets the model output (at a certain training iteration), and finally, the fourth row presents the same output with increased contrast for extra clarity.
(a) Iter = 2k
(b) Iter = 6k
(c) Iter = 8k
(d) Iter = 10k
(e) Iter = 13k
(f) Iter = 50k
Figure 22: Expanded staircase examples for the cycles 16 task. The first row and second row represent the input image and the final ground truth CoS frame respectively. The third image represents the output of the model and the fourth image is similar to the third image with a higher contrast.
(a) Iter = 3k
(b) Iter = 5k
(c) Iter = 6k
(d) Iter = 8k
(e) Iter = 11k
(f) Iter = 50k
Figure 23: Additional staircase example for the maze (rectangular) task. The first row and second row show the input image and the final ground truth CoS frame respectively. The third image represents the output of the model and the fourth image is similar to the third image with a higher contrast.
F.3 CoS generations of the models
All of our CoS models that were large enough to learn the tasks were also able to generate the CoS frame(s) with a rather high quality. For reference, here we report some of the sketches that our inductive CoS model generated for different tasks in Figures 24 to 28.
Figure 24: Generated CoS frames by our inductive CoS model for an example from the cycles task. The input is shown on the left, followed by the generated frames from left to right.
Figure 25: Generated CoS frames by our inductive CoS model for a strings task example. Input is shown on the left, followed by the generated frames are shown from left to right.
Figure 26: Generated CoS frames by our inductive CoS model for a maze (rect.) task example. Input is shown on the left, followed by the generated frames are shown from left to right.
Figure 27: Generated CoS frames by our inductive CoS model for a maze (circ.) task example. Input is shown on the left, followed by the generated frames are shown from left to right.
Figure 28: Generated CoS frames by our inductive CoS model for a PVR task example. Input is shown on the left, followed by the generated frames are shown from left to right.